OpenStudy (anonymous):

(-4+x) -12 What is the value of the expression when x=6?

OpenStudy (anonymous):

Can't be.

OpenStudy (anonymous):

just plug in the 6 where the x is and solve the equation

OpenStudy (whovianchick):

(-4+6)-12 (2)-12 -10 Negative ten

OpenStudy (anonymous):

(-4+6) -12 2-12 -10

OpenStudy (anonymous):

\[\left( -4+\left( 6 \right) \right)-12=?\]

OpenStudy (shreenithaa):

ooh sorry i thought the -12 was =12 sorry

OpenStudy (whovianchick):

@shreenithaa No the sign before the twelve is a subtraction sign, not an equal sign

OpenStudy (whovianchick):

There you go

OpenStudy (anonymous):

Its fine!

OpenStudy (anonymous):

Thanks guys!"

OpenStudy (whovianchick):

:) no prob